# The Obstacle Problem Arising from the American Chooser Option

### v2

@@ 3.1. Existence and uniqueness in a bounded region
Define a penalty function $\beta_{\varepsilon}(t)\in C^{\infty}(\mathbb{R})$ with $\varepsilon>0$ satisfying
(3.2) $\begin{cases}\ \beta_{\varepsilon}(t)\leqslant 0,\ \beta^{\prime}_{\varepsilon}(t)\geqslant 0\ \ \text{and}\ \ \beta^{\prime\prime}_{\varepsilon}(t)\leqslant 0\ \ \text{for \ all}\ \ t\in\mathbb{R},\\ \ \beta_{\varepsilon}(t)=0\ \ \text{if}\ \ t\geqslant\varepsilon,\ \beta_{\varepsilon}(0)=-\mathcal{K}_{0}\ \ \text{where}\ \ \mathcal{K}_{0}:=2\{(q+r)e^{n}+2rK_{c}+5r\},\\ \ \displaystyle{\lim_{\varepsilon\to 0}\beta_{\varepsilon}(t)}=0\ \ \text{if}\ \ t>0,\quad\displaystyle{\lim_{\varepsilon\to 0}\beta_{\varepsilon}(t)}=-\infty\ \ \text{if}\ \ t<0\end{cases}$
and $\varphi_{\varepsilon}(t)\in C^{{\infty}}(\mathbb{R})$ satisfying (2.1). We then consider the following penalized problem;
@@
For each fixed $n\in\mathbb{N}\setminus\{0\}$ with $n>\textup{max}\{|{x_{c}}(0)|,|{x_{p}}(0)|\}$, there exists a unique solution $V_{n,\varepsilon}\in W^{1,2}_{p}(\Omega_{T}^{n})\cap C(\overline{\Omega_{T}^{n}})$ to the problem (3.3), where $1<p<\infty.$
@@
We apply the Schauder fixed point theorem [10, 280p] to the following setting. Set $\mathcal{B}:=C(\overline{\Omega_{T}^{n}})$ and $\mathcal{D}:=\{w\in\mathcal{B}:w\geqslant 0\}$. Then $\mathcal{D}$ is a closed convex subset of the Banach space $\mathcal{B}.$ For each $w\in\mathcal{D}$, let $u\in W^{1,2}_{p}(\Omega_{T}^{n})$ be the unique solution to
@@
Note that the well-posedness of the solution to (3.5) in $W^{1,2}_{p}$ can be found in [19], and the relevant $W^{1,2}_{p}(\Omega_{T}^{n})$ estimate implies
$\|u\|_{W^{1,2}_{p}(\Omega_{T}^{n})}\leqslant\mathcal{K}(\|e^{n}+e^{-n}\|_{W^{1}_{p}([0,T))}+\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(w-J_{\varepsilon})\|_{L^{p}(\Omega_{T}^{n})})$
for some constant $\mathcal{K}>0$. Moreover, the parabolic Sobolev embedding theorem [19] yields that $u$ is Hölder continuous in $\overline{\Omega_{T}^{n}}$. Thus $u\in\mathcal{B}.$ Now we define operator $\mathcal{F}:\mathcal{D}\rightarrow\mathcal{B}$ by $\mathcal{F}(w):=u$. In order to use the Schauder fixed point theorem in $\mathcal{B},$ it suffices to show the following three properties:
@@
    $\mathcal{F}(\mathcal{D})\subset\mathcal{D}$;
@@
    $\mathcal{F}(\mathcal{D})$ is precompact in $\mathcal{B}$.
We shall prove it in Appendix. Utilizing the Schauder fixed point theorem, we obtain the solution $V_{n,\varepsilon}$ of the problem (3.3). In particular, $V_{n,\varepsilon}\in W^{1,2}_{p}(\Omega^{n}_{T})\cap C(\overline{\Omega^{n}_{T}})$ for each $1<p<\infty.$
Once we have obtained the solution $V_{n,\varepsilon}$ of the problem (3.3), we focus on the convergence of $V_{n,\varepsilon}$ as $\varepsilon\rightarrow 0^{+}$. Our claim is that the limit exists and is in fact, a solution to the obstacle problem (3.1).
@@
Let $V_{n,\varepsilon}$ be the solution of (3.1). Then for each $n\in\mathbb{N},$ there exists some constant $\mathcal{K}>0$ independent of $\varepsilon\in(0,1)$ such that
@@
Next, observe that $\partial_{x}J_{\varepsilon}=(1-\varphi^{\prime}_{\varepsilon}(C-P))\partial_{x}P+\varphi^{\prime}_{\varepsilon}(C-P)\partial_{x}C$ and $0\leqslant\varphi^{\prime}_{\varepsilon}\leqslant 1$. Combining these estimates, we deduce
(3.6) $-e^{x}\leqslant\partial_{x}J_{\varepsilon}\leqslant e^{x}\quad\text{in}\ \ \Omega^{n}_{T}.$
@@
[…]
By direct computation together with the choice of $\mathcal{K}_{0}$, $J_{\varepsilon}$ satisfies that in $\Omega^{n}_{T}$,
$\partial_{\tau}J_{\varepsilon}-\mathcal{L}J_{\varepsilon}$ $=\varphi^{\prime}_{\varepsilon}(C-P)\big{[}\partial_{\tau}C-\mathcal{L}C\big{]}+\big{\{}1-\varphi_{\varepsilon}^{\prime}(C-P)\big{\}}\big{[}\partial_{\tau}P-\mathcal{L}P\big{]}$
$\ \ -\frac{\sigma^{2}}{2}\varphi^{\prime\prime}_{\varepsilon}(C-P)\big{[}\partial_{x}(C-P)\big{]}^{2}-r\varphi^{\prime}(C-P)\big{[}C-P\big{]}+r\varphi_{\varepsilon}(C-P)$
$\leqslant\ \varphi^{\prime}(\cdot)\big{|}qe^{n}-rK_{c}\big{|}+\big{\{}1-\varphi^{\prime}(\cdot)\big{\}}\big{|}-qe^{n}+rK_{p}\big{|}+2r\big{[}\max\limits_{\Omega^{n}_{T}}\big{|}C-P\big{|}+1\big{]}$
@@
for sufficiently small $10<\varepsilon<$ 1. Consequently, we deduce
$\begin{cases}\partial_{\tau}J_{\varepsilon}-\mathcal{L}J_{\varepsilon}+\beta_{\varepsilon}(0)\leqslant 0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}J_{\varepsilon}(\tau,-n)\geqslant-e^{-n}=\partial_{x}V_{n,\varepsilon}(\tau,-n)\quad\text{and}\quad\partial_{x}J_{\varepsilon}(\tau,n)\leqslant e^{n}=\partial_{x}V_{n,\varepsilon}(\tau,n)\quad\text{for }\ \tau\in[0,T],\\ J_{\varepsilon}(0,x)=V_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
Hence, $J_{\varepsilon}$ is a subsolution to (3.3) and by the comparison principle [6, 53p]
@@
$-\mathcal{K}_{0}=\beta_{\varepsilon}(0)\leqslant\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\leqslant 0.$
Next, our claim is that $V_{n,\varepsilon}$ is bounded. To this end, define $h(\tau,x):=e^{x}+e^{-n}x^{2}+\mathcal{K}_{1}$ where $\mathcal{K}_{1}$ is a constant satisfying $\mathcal{K}_{1}\geqslant\max\{(1-\frac{q}{r}-\frac{\sigma^{2}}{2r})^{2}+\frac{\sigma^{2}}{r},K_{p}+6\}.$ This implies $h\geqslant J_{\varepsilon}+\varepsilon$ and hence $\beta_{\varepsilon}(h-J_{\varepsilon})=0$ in $\Omega_{T}^{n}.$ Moreover in $\Omega_{T}^{n}$,
$\partial_{\tau}h-\mathcal{L}h=qe^{x}+e^{-n}\{rx^{2}-(2r-2q-\sigma^{2})x-\sigma^{2}\}+r\mathcal{K}_{1}\geqslant 0.$
Thus, $h$ satisfies
$\begin{cases}\partial_{\tau}h-\mathcal{L}h+\beta_{\varepsilon}(h-J_{\varepsilon})\geqslant 0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}h(\tau,-n)=(1-2n)e^{-n}\leqslant\partial_{x}V_{n,\varepsilon}(\tau,-n)\quad\text{and}\\ \partial_{x}h(\tau,n)=2ne^{-n}+e^{n}\geqslant\partial_{x}V_{n,\varepsilon}(\tau,n)\quad\text{for }\ \tau\in[0,T],\\ h(0,x)\geqslant V_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
By the comparison principle [6, 53p],
$V_{n,\varepsilon}\leqslant h\leqslant e^{n}+e^{-n}n^{2}+\mathcal{K}_{1}\quad\text{in}\ \ \Omega^{n}_{T}.\qed$
[…]
@@
For each fixed $n\in\mathbb{N}\setminus\{0\}$ with $n>\textup{max}\{|{x_{c}}(0)|,|{x_{p}}(0)|\}$, there exists a unique solution $V_{n}\in C(\overline{\Omega^{n}_{T}})\cap W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))$ to the problem (3.1), where $1<p<\infty$ and $\varrho>0$.
@@
For the solution $V_{n,\varepsilon}$ to we consider the $W^{1,2}_{p}$ estimate in the domain $\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}),$ where $1<p<\infty$ and $\varrho>0$ and $C^{\frac{\alpha}{2},\alpha}$ estimate for $V_{n,\varepsilon}$ [19, 14]:
$\|V_{n,\varepsilon}\|_{W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))}$ $\leqslant\mathcal{K}\Big{(}\|V_{n,\varepsilon}\|_{L^{\infty}(\Omega^{n}_{T})}++\|\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\|_{L^{p}(\Omega^{n}_{T})}+\|e^{-n}+e^{n}\|_{W^{1}_{p}([0,T))}$
$\quad\quad+\|\varphi_{\epsilon}(C-P)+P\|_{W^{2}_{p}((-n,-n)\setminus(\bar{x}-\delta/2,\bar{x}+\delta/2))}\Big{)}$
$\leqslant\mathcal{K}^{\prime},$
and
$\|V_{n,\varepsilon}\|_{C^{\frac{\alpha}{2},\alpha}(\overline{\Omega^{n}_{T}})}\leqslant\mathcal{K}^{\prime\prime}.$
Note that the constants $\mathcal{K}^{\prime}$ and $\mathcal{K}^{\prime\prime}$ are independent of $\varepsilon\in(0,\delta/2)$ due to Lemma 3.2 and the property of $\varphi_{\varepsilon}$ in (2.1). Therefore, for each large $n\in\mathbb{N}$, $V_{n,\varepsilon}$ is bounded in $C(\overline{\Omega^{n}_{T}})\cap W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,x))$ for $\varepsilon\in(0,\delta/2)$, and hence there exists $V_{n}\in C(\overline{\Omega^{n}_{T}})\cap W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))$ for all $1<p<\infty$ and $\varrho>0$ such that
$V_{n,\varepsilon}\rightarrow V_{n}\quad$ $\text{in}\quad C(\overline{\Omega^{n}_{T}}),$
@@
up to a subsequence, as $\varepsilon\rightarrow 0^{+}$. Moreover, the Sobolev embedding theorem yields
$\partial_{x}V_{n,\varepsilon}\rightarrow\partial_{x}V_{n}\quad$ $\text{in}\quad C(\overline{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})}).$
Now, let us verify that $V_{n}$ is a solution to (3.1). Fix any $\psi\in C_{c}^{\infty}(\{V_{n}>J\}\cap\Omega^{n}_{T})$. Then there exist small $\varrho,\delta>0$ such that
$\text{supp}(\psi)\subset\{V_{n}>J\}\cap\Omega^{n}_{T}\setminus B_{\varrho}(0,x_{0})\quad\text{and}\quad\min\limits_{\text{supp}(\psi)}(V_{n}-J)>2\delta.$
@@
Then for such small $\varepsilon>0$
$V_{n,\varepsilon}-J_{\varepsilon}>\delta>\varepsilon,$
and this implies
$\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=0\ \ \text{in}\ \text{supp}(\psi).$
Therefore, we obtain
[…]
Thus, by (3.7), taking $\varepsilon\rightarrow 0^{+}$ yields $\partial_{\tau}V_{n}-\mathcal{L}V_{n}=0$ a.e. in $\{V_{n}>J\}\cap\Omega^{n}_{T}.$ We next show that $V_{n}$ is a supersolution in $\Omega^{n}_{T}$. Fix any non-negative $\psi\in C_{c}^{\infty}(\Omega^{n}_{T})$. Then Lemma 3.2 implies that for all $\varepsilon\in(0,1),$
[…]
We refer to (3.7) and hence taking $\varepsilon\rightarrow 0^{+}$ yields
[…]
Therefore, $V_{n}$ is a solution to (3.1).
Now, we prove the uniqueness of $V_{n}.$ Let $V_{n}$ and $\tilde{V}_{n}$ be two solutions of (3.1). Then, define $\mathcal{N}:=\{(\tau,x)\in\overline{\Omega^{n}_{T}}:{\tilde{V}_{n}}(\tau,x)>V_{n}(\tau,x)\}$ and suppose that $\mathcal{N}$ is nonempty. Notice that $V_{n},\tilde{V}_{n}$ are continuous functions. Hence the set $\mathcal{N}$ is open, contains a ball inside. In the set $\mathcal{N},$ we have $\tilde{V}_{n}>J$ so
$\partial_{\tau}\tilde{V}_{n}-\mathcal{L}\tilde{V}_{n}=0\ \ \text{in}\ \ \mathcal{N}.$
Next, we define an operator $\mathcal{M}$ by
$\mathcal{M}u:=-u+\beta\cdot Du$
where $\beta$ is the inward pointing unit vector field with its $\tau$-direction $\beta^{\tau}=0$ on $\mathcal{P}\mathcal{N}$. Since both $V_{n},\tilde{V}_{n}$ and each of their partial derivative with respect to $x$ are continuous by the Sobolev embedding, $\mathcal{M}(\tilde{V}_{n}-V_{n})=0$ on the parabolic boundary $\mathcal{P}\mathcal{N}$. Thus,
$\begin{cases}\partial_{\tau}{(\tilde{V}_{n}-V_{n})}-\mathcal{L}{(\tilde{V}_{n}-V_{n})}\leqslant 0\quad&\text{in }\ \mathcal{N},\\ \mathcal{M}{(\tilde{V}_{n}-V_{n})}=0\quad&\text{on}\ \ \mathcal{P}\mathcal{N}.\end{cases}$
By the comparison principle [14, Corollary 7.4], it follows that $\tilde{V}_{n}-V_{n}\leqslant 0$ and this contradicts the definition of $\mathcal{N}$ and hence $\mathcal{N}=\emptyset$. Applying the same procedure as above to $V_{n}-\tilde{V}_{n}$ in the set $\{V_{n}>\tilde{V}_{n}\}$ gives $\{V_{n}>\tilde{V}_{n}\}=\emptyset$. This proves $\tilde{V}_{n}\equiv V_{n}$ in $\Omega_{T}^{n}.$

### v3

@@ 3.1. Existence and uniqueness in a bounded region
Define a penalty function $\beta_{\varepsilon}(\cdot)\in C^{\infty}(\mathbb{R})$ with $\varepsilon>0$ satisfying
(3.2) $\begin{cases}\beta_{\varepsilon}(\lambda)\leqslant 0,\ \beta^{\prime}_{\varepsilon}(\lambda)\geqslant 0\ \ \text{and}\ \ \beta^{\prime\prime}_{\varepsilon}(\lambda)\leqslant 0\ \ \text{for \ all}\ \ \lambda\in\mathbb{R},\\ \beta_{\varepsilon}(\lambda)=0\ \ \text{if}\ \ \lambda\geqslant\varepsilon,\ \beta_{\varepsilon}(0)=-\mathcal{K}_{0}\ \ \text{where}\ \ \mathcal{K}_{0}:=2\{(q+r)e^{n}+2rK_{c}+5r\},\\ \displaystyle{\lim_{\varepsilon\to 0}\beta_{\varepsilon}(\lambda)}=0\ \ \text{if}\ \ \lambda>0,\quad\displaystyle{\lim_{\varepsilon\to 0}\beta_{\varepsilon}(\lambda)}=-\infty\ \ \text{if}\ \ \lambda<0\end{cases}$
and $\varphi_{\varepsilon}(\cdot)\in C^{{\infty}}(\mathbb{R})$ satisfying (2.1). We then consider the following penalized problem;
@@
For each fixed $n\in\mathbb{N}\setminus\{0\}$ with $n>\max\{|{x_{c}}(0)|,|{x_{p}}(0)|\}$, there exists a solution $V_{n,\varepsilon}\in W^{1,2}_{p}(\Omega_{T}^{n})\cap C(\overline{\Omega_{T}^{n}})$ to the problem (3.3), where $1<p<\infty.$
@@
We apply the Schauder fixed point theorem [11, 280p] to the following setting. Set $\mathcal{X}:=C(\overline{\Omega_{T}^{n}})$ and $\mathcal{A}:=\{w\in\mathcal{X}:w\geqslant 0\}$. Then $\mathcal{A}$ is a closed convex subset of the Banach space $\mathcal{X}.\$For each $w\in\mathcal{A}$, let $u\in W^{1,2}_{p}(\Omega_{T}^{n})$ be the solution to
@@
Note that the well-posedness of the solution to (3.5) and the relevant estimate in $W^{1,2}_{p}(\Omega_{T}^{n})$ for $1<p<\infty$ can be found in [15, Theorem 9.1, p.341]:
$\|u\|_{W^{1,2}_{p}(\Omega_{T}^{n})}\leqslant\mathcal{K}\left(T(e^{n}+e^{-n})+\|J_{\varepsilon}\|_{W^{2}_{p}((-n,n))}+\|\beta_{\varepsilon}(w-J_{\varepsilon})\|_{L^{p}(\Omega_{T}^{n})}\right)$
for some constant $\mathcal{K}>0$. Moreover, the parabolic Sobolev embedding theorem [15, Chapter II. Lemma 3.3] yields that $u$ is Hölder continuous in $\overline{\Omega_{T}^{n}}$. Thus $u\in\mathcal{X}.$ Now we define an operator $\mathcal{F}:\mathcal{A}\rightarrow\mathcal{X}$ by $\mathcal{F}(w):=u$. In order to use the Schauder fixed point theorem in $\mathcal{X},$ it suffices to show the following three properties:
@@
    $\mathcal{F}(\mathcal{A})\subset\mathcal{A}$;
@@
    $\mathcal{F}(\mathcal{A})$ is precompact in $\mathcal{X}$.
We shall prove it in Appendix D. Utilizing the Schauder fixed point theorem, we obtain the solution $V_{n,\varepsilon}$ of the problem (3.3). In particular, $V_{n,\varepsilon}\in W^{1,2}_{p}(\Omega^{n}_{T})\cap C(\overline{\Omega^{n}_{T}})$ for each $1<p<\infty.$
Next, we investigate the convergence of $V_{n,\varepsilon}$ as $\varepsilon\to 0^{+}$. We establish that the limit exists and is a solution to the obstacle problem (3.1).
@@
Let $V_{n,\varepsilon}$ be the solution of (3.3). Then for each $n\in\mathbb{N},$ there exists some constant $\mathcal{K}=\mathcal{K}(n)>0$ independent of $\varepsilon\in(0,1)$ such that
@@
Next, observe that $\partial_{x}J_{\varepsilon}=\{1-\varphi^{\prime}_{\varepsilon}(C-P)\}\partial_{x}P+\varphi^{\prime}_{\varepsilon}(C-P)\partial_{x}C$ and $0\leqslant\varphi^{\prime}_{\varepsilon}(\cdot)\leqslant 1$. Combining these estimates, we deduce
(3.6) $e^{x}\leqslant\partial_{x}J_{\varepsilon}\leqslant e^{x}\quad\text{in}\ \ \Omega^{n}_{T}.$
@@
$\partial_{\tau}C-\mathcal{L}C=(qe^{x}-rK_{c})\mathbf{1}_{\{C=e^{x}-K_{c}\}}\quad\text{and}\quad\partial_{\tau}P-\mathcal{L}P=(-qe^{x}+rK_{p})\mathbf{1}_{\{P=K_{p}-e^{x}\}}.$
Direct computation and the choice of $\mathcal{K}_{0}$ yield the following in $\Omega_{T}^{n}$:
$\partial_{\tau}J_{\varepsilon}-\mathcal{L}J_{\varepsilon}$ $=\varphi^{\prime}_{\varepsilon}(C-P)\left(\partial_{\tau}C-\mathcal{L}C\right)+\left(1-\varphi^{\prime}_{\varepsilon}(C-P)\right)\left(\partial_{\tau}P-\mathcal{L}P\right)$
$\quad-\frac{\sigma^{2}}{2}\varphi^{\prime\prime}_{\varepsilon}(C-P)\left[\partial_{x}(C-P)\right]^{2}-r\varphi^{\prime}_{\varepsilon}(C-P)(C-P)+r\varphi_{\varepsilon}(C-P)$
$\leqslant\varphi^{\prime}_{\varepsilon}(\cdot)|qe^{n}-rK_{c}|+(1-\varphi^{\prime}_{\varepsilon}(\cdot))|-qe^{n}+rK_{p}|$
$\quad+2r\left(\max_{\Omega_{T}^{n}}|C-P|+1\right)$
@@
for sufficiently small $0<\varepsilon<1$. Consequently, we deduce
$\begin{cases}\partial_{\tau}J_{\varepsilon}-\mathcal{L}J_{\varepsilon}+\beta_{\varepsilon}(0)\leqslant 0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}J_{\varepsilon}(\tau,-n)\geqslant-e^{-n}=\partial_{x}V_{n,\varepsilon}(\tau,-n)\quad\text{and}\\ \partial_{x}J_{\varepsilon}(\tau,n)\leqslant e^{n}=\partial_{x}V_{n,\varepsilon}(\tau,n)\quad\text{for }\ \tau\in[0,T],\\ J_{\varepsilon}(0,x)=V_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
Hence, $J_{\varepsilon}$ is a subsolution to (3.3) and by the comparison principle [8, Theorem 17, p.53]
@@
$\mathcal{K}_{0}=\beta_{\varepsilon}(0)\leqslant\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\leqslant 0.$
We next show that $V_{n,\varepsilon}$ is bounded from above. To this end, define
$\overline{V}(\tau,x):=e^{x}+e^{-n}x^{2}+\mathcal{K}_{1},$
where $\mathcal{K}_{1}$ is a constant satisfying
$\mathcal{K}_{1}\geqslant\max\left\{\bigg(1-\frac{q}{r}-\frac{\sigma^{2}}{2r}\right)^{2}+\frac{\sigma^{2}}{r},\;K_{p}+6\bigg\}.$
This implies $\overline{V}\geqslant J_{\varepsilon}+\varepsilon$ and hence $\beta_{\varepsilon}(\overline{V}-J_{\varepsilon})=0$ in $\Omega_{T}^{n}.$ Moreover in $\Omega_{T}^{n}$,
$\partial_{\tau}\overline{V}-\mathcal{L}\overline{V}=qe^{x}+e^{-n}\{rx^{2}-(2r-2q-\sigma^{2})x-\sigma^{2}\}+r\mathcal{K}_{1}\geqslant 0.$
Thus, $\overline{V}$ satisfies
$\begin{cases}\partial_{\tau}\overline{V}-\mathcal{L}\overline{V}+\beta_{\varepsilon}(\overline{V}-J_{\varepsilon})\geqslant 0\quad\text{in }\ \Omega^{n}_{T},\\ \partial_{x}\overline{V}(\tau,-n)=(1-2n)e^{-n}\leqslant\partial_{x}V_{n,\varepsilon}(\tau,-n)\quad\text{and}\\ \partial_{x}\overline{V}(\tau,n)=2ne^{-n}+e^{n}\geqslant\partial_{x}V_{n,\varepsilon}(\tau,n)\quad\text{for }\ \tau\in[0,T],\\ \overline{V}(0,x)\geqslant V_{n,\varepsilon}(0,x)\quad\text{for }\ x\in(-n,n).\end{cases}$
By the comparison principle [8, Theorem 17, p.53],
$V_{n,\varepsilon}\leqslant\overline{V}\leqslant e^{n}+e^{-n}n^{2}+\mathcal{K}_{1}\quad\text{in}\ \ \Omega^{n}_{T}.\qed$
We now establish the existence of the solution $V_{n}$ to problem (3.1). To this end, let us recall the unique point $\bar{x}$ satisfying $C(0,\bar{x})=P(0,\bar{x})$ from (2.9). For each $\varrho>0$, we define the ball $B_{\varrho}(0,\bar{x}):=\{(\tau,x)\in\mathbb{R}^{2}:|(\tau,x)-(0,\bar{x})|<\varrho\}$.
@@
For each fixed $n\in\mathbb{N}$ with $n>\max\{|{x_{c}}(0)|,|{x_{p}}(0)|\}$ and $\varrho>0$, there exists a solution $V_{n}\in W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))$ to the problem (3.1), where $1<p<\infty$. Moreover, $V_{n}\in C(\overline{\Omega^{n}_{T}})$.
@@
Since $J(0,x)=\max\{C(0,x),P(0,x)\}$ is not differentiable at the point $(0,\bar{x})$, we consider the following $W^{1,2}_{p}$-estimate in the domain $\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})$ for $1<p<\infty$ and $\varrho>0$:
$\|V_{n,\varepsilon}\|_{W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))}$ $\leqslant\mathcal{K}\Big(\|V_{n,\varepsilon}\|_{L^{\infty}(\Omega^{n}_{T})}+\|\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\|_{L^{\infty}(\Omega^{n}_{T})}+T(e^{n}+e^{-n})$
$\quad\quad+\|\varphi_{\varepsilon}(C-P)+P\|_{W^{2}_{p}((-n,n)\setminus(\bar{x}-\varrho/2,\bar{x}+\varrho/2))}\Big)$
$\leqslant\mathcal{K}^{\prime}.$
By the Sobolev embedding theorem, we deduce that for every $(\tau_{0},x_{0})\in\overline{\Omega_{T}^{n}}\setminus\{(0,\bar{x})\}$, $V_{n,\varepsilon}$ is $C^{\frac{\alpha}{2},\alpha}$ in some neighborhood of $(\tau_{0},x_{0})$ in $\overline{\Omega_{T}^{n}}$. Note that the constant $\mathcal{K}^{\prime}$ is independent of $\varepsilon\in(0,\varrho/2)$ due to Lemma 3.2 and the property of $\varphi_{\varepsilon}$ in (2.1). Therefore, for each large $n\in\mathbb{N}$, $V_{n,\varepsilon}$ is bounded in $W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))\cap C(\overline{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})})$ for $\varepsilon\in(0,\varrho/2)$, and hence there exists $V_{n}\in W^{1,2}_{p}(\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x}))\cap C(\overline{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})})$ for all $1<p<\infty$ and $\varrho>0$ such that
@@
(3.8) $V_{n,\varepsilon}\rightarrow V_{n}\quad$ $\text{in}\quad C(\overline{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})}),$
up to a subsequence, as $\varepsilon\rightarrow 0^{+}$.
Now, we prove that $V_{n}$ satisfies (3.1). Observe from the inequality $V_{n,\varepsilon}\geqslant J_{\varepsilon}$, the uniform convergence $V_{n,\varepsilon}\to V_{n}$ and $J_{\varepsilon}\to J$ in $\Omega_{T}^{n}$, as $\varepsilon\to 0^{+}$, yields $V_{n}\geqslant J$ in $\Omega_{T}^{n}$. Fix any $\psi\in C_{c}^{\infty}(\{V_{n}>J\}\cap\Omega^{n}_{T})$. Then, there exist small $\varrho,\delta>0$ such that
$\text{supp}(\psi)\subset\{V_{n}>J\}\cap\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})\quad\text{and}\quad\min\limits_{\text{supp}(\psi)}(V_{n}-J)>2\delta.$
@@
Then, for such small $\varepsilon>0$, it follows that $V_{n,\varepsilon}-J_{\varepsilon}>\delta>\varepsilon$, and this implies $\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})=0$ in $\text{supp}(\psi)$. Consequently,
$\iint_{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})}\big(\partial_{\tau}V_{n,\varepsilon}-\mathcal{L}V_{n,\varepsilon}\big)\psi\,d\tau\,dx=0.$
By the weak convergence (3.7), taking $\varepsilon\rightarrow 0^{+}$ yields
$\iint_{\Omega^{n}_{T}\setminus B_{\varrho}(0,\bar{x})}\big(\partial_{\tau}V_{n}-\mathcal{L}V_{n}\big)\psi\,d\tau\,dx=0.$
Since $\psi\in C_{c}^{\infty}(\{V_{n}>J\}\cap\Omega^{n}_{T})$ is arbitrary, we deduce that $\partial_{\tau}V_{n}-\mathcal{L}V_{n}=0$ a.e. in $\{V_{n}>J\}\cap\Omega^{n}_{T}.$
To prove $\partial_{\tau}V_{n}-\mathcal{L}V_{n}\geqslant 0$ a.e. in $\Omega^{n}_{T}$, let us fix a non-negative $\psi\in C_{c}^{\infty}(\Omega^{n}_{T})$. Then since $\beta_{\varepsilon}(\cdot)\leqslant 0$, $V_{n,\varepsilon}$ satisfies that for all small $\varepsilon>0$,
$\iint_{\Omega^{n}_{T}}({\partial_{\tau}V_{n,\varepsilon}-\mathcal{L}V_{n,\varepsilon}})\psi\,d\tau\,dx=\iint_{\Omega^{n}_{T}}-\beta_{\varepsilon}(V_{n,\varepsilon}-J_{\varepsilon})\psi\,d\tau\,dx\geqslant 0.$
By the weak convergence (3.7), taking $\varepsilon\rightarrow 0^{+}$ yields
$\iint_{\Omega^{n}_{T}}({\partial_{\tau}V_{n}-\mathcal{L}V_{n}})\psi\,d\tau\,dx\geqslant 0.$
Since $\psi$ is an arbitrary non-negative test function in $C_{c}^{\infty}(\Omega_{T}^{n})$, we deduce that $\partial_{\tau}V_{n}-\mathcal{L}V_{n}\geqslant 0$ a.e. in $\Omega_{T}^{n}$. Combining all, we conclude that $V_{n}$ is a solution of (3.1).
Finally, by the convergence (3.8) for each $\varrho>0$ and the continuity of the initial datum $V_{n}(0,x)=\max\{C(0,x),P(0,x)\}$, we obtain that $V_{n}\in C(\overline{\Omega_{T}^{n}})$.